OpenStudy (anonymous):

Solve for x: −4|x + 5| = −16

OpenStudy (alexandervonhumboldt2):

first divide both sides by -4. we will get |x + 5| = 4 here is now to solve absolute value equations: x+5=4 => x=-1 -x-5=4 => x=-9 thus the roots are -1 and -9